# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 5 Running Examples Continued
$C\langle E,F\rangle=\{\,\{\,\#\,\},\;\newline \{\,\$,\;f(\$,\$),\;f(\$,f(\$,\$)),\;f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$))\,\},\;\newline \{\,f(\#,\#),\;f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;f(f(\#,\#),f(\#,\#))\,\},\newline \;\{\,\mathsterling,\;f(\#,\$),\;f(\#,\mathsterling),\;f(\$,\#),\;f(\$,\mathsterling),\;f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}\,\}$,
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;f(\#/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\#/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\$/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\$/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\;f(\$/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\mathsterling/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\#/_{\Theta\langle E,F\rangle},f(\#,\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;\newline f(f(\#,\#)/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;\,f(f(\#,\#)/_{\Theta\langle E,F\rangle},f(\#,\#))/_{\Theta\langle E,F\rangle}\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle}\;\;\newline f(\$/_{\Theta\langle E,F\rangle},f(\#,\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;f(\mathsterling/_{\Theta\langle E,F\rangle},f(\#,\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(f(\#,\#)/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;f(f(\#,\#)/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle}\,\}$.
$C\langle F,E\rangle=\{\,\{\,\$\,\},\;\newline \{\,\#,\;f(\#,\#),\,f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;f(f(\#,\#),f(\#,\#))\,\},\;\newline \{\,f(\$,\$),\;f(\$,f(\$,\$)),\;f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$))\,\},\newline \;\{\,\mathsterling,\;f(\$,\#),\;f(\$,\mathsterling),\;f(\#,\$),\;f(\#,\mathsterling),\;f(\mathsterling,\$),\;f(\mathsterling,\#),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling)\,\}\,\}$,
$R\langle F,E\rangle=\{\,\$\dot{\rightarrow}\$/_{\Theta\langle F,E\rangle},\;\;\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\$/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;\;f(\$/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\$/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\#/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\;f(\#/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\mathsterling/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\$/_{\Theta\langle F,E\rangle},f(\$,\$)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;\;\newline f(f(\$,\$)/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;\;\,f(f(\$,\$)/_{\Theta\langle F,E\rangle},f(\$,\$))/_{\Theta\langle F,E\rangle}\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle}\;\;\newline f(\#/_{\Theta\langle F,E\rangle},f(\$,\$)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;\;f(\mathsterling/_{\Theta\langle F,E\rangle},f(\$,\$)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(f(\$,\$)/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle},\;\;f(f(\$,\$)/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle}\,\}$.
$C\langle E\cup F,\emptyset\rangle=\{\,\{\,\#,\;\$,\;\mathsterling,\;\newline f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;f(f(\#,\#),f(\#,\#)),\;\;\newline f(\$,f(\$,\$)),\;f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$)),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}\,\}$,
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathsterling\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
@@
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathsterling\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\flat\dot{\rightarrow}\flat/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},\flat/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;f(\flat/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#\,\}$.
$C\langle F,E\rangle=\{\,\#/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle}\,\}$,
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle F,E\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\flat\dot{\rightarrow}\mathsterling_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\newline f(\$/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\$/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\$/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\newline f(\mathsterling/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\mathsterling/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\mathsterling/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle}\,\}$.
@@
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathsterling\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\flat\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
@@
$C\langle E,F\rangle=\{\,\#/_{\Theta\langle E,F\rangle},\;\$/_{\Theta\langle E,F\rangle}\,\}$.
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;\newline f(\$/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle},\;f(\$/_{\Theta\langle E,F\rangle},\$/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\$/_{\Theta\langle E,F\rangle}\;\}$.
@@
$C\langle F,E\rangle=\{\,\{\,\#,\;f(\#,\#)\,\},\;\{\,\$\,\},\;\{\,f(\#,\$)\,\},\;\{\,f(\$,\#)\,\},\;\{\,f(\$,\$)\,\}\,\}$.
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\$\dot{\rightarrow}\$/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},f(\#/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#,\$)/_{\Theta\langle F,E\rangle},\;\newline f(\$/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\#)/_{\Theta\langle F,E\rangle},\;f(\$/_{\Theta\langle F,E\rangle},\$/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\$,\$)/_{\Theta\langle F,E\rangle}\,\}$.
@@
$C\langle E\cup F,\emptyset\rangle=\{\,\{\,\#,\;\$,\;f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}\,\}$.
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\$\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,E\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.

### v2

@@ 5 Running Examples Continued
[…]
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;f(\#/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\#/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\mathdollar/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\mathdollar/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\;f(\mathdollar/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\mathsterling/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\#/_{\Theta\langle E,F\rangle},f(\#,\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;\newline f(f(\#,\#)/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle},\;\;\,f(f(\#,\#)/_{\Theta\langle E,F\rangle},f(\#,\#))/_{\Theta\langle E,F\rangle}\dot{\rightarrow}f(\#,\#)/_{\Theta\langle E,F\rangle}\;\;\newline f(\mathdollar/_{\Theta\langle E,F\rangle},f(\#,\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(\mathsterling/_{\Theta\langle E,F\rangle},f(\#,\#)/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;\newline f(f(\#,\#)/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle},\;\;f(f(\#,\#)/_{\Theta\langle E,F\rangle},\mathsterling/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle E,F\rangle}\,\}$.
$C\langle F,E\rangle=\{\,\{\,\mathdollar\,\},\;\newline \{\,\#,\;f(\#,\#),\,f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;f(f(\#,\#),f(\#,\#))\,\},\;\newline \{\,f(\mathdollar,\mathdollar),\;f(\mathdollar,f(\mathdollar,\mathdollar)),\;f(f(\mathdollar,\mathdollar),\mathdollar),\;f(f(\mathdollar,\mathdollar),f(\mathdollar,\mathdollar))\,\},\newline \;\{\,\mathsterling,\;f(\mathdollar,\#),\;f(\mathdollar,\mathsterling),\;f(\#,\mathdollar),\;f(\#,\mathsterling),\;f(\mathsterling,\mathdollar),\;f(\mathsterling,\#),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\mathdollar,\mathdollar)),\;f(\mathsterling,f(\mathdollar,\mathdollar)),\;f(f(\mathdollar,\mathdollar),\#),\;f(f(\mathdollar,\mathdollar),\mathsterling),\;\newline f(\mathdollar,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\mathdollar),\;f(f(\#,\#),\mathsterling)\,\}\,\}$,
$R\langle F,E\rangle=\{\,\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle F,E\rangle},\;\;\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\mathdollar/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\;\;f(\mathdollar/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\mathdollar/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\#/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\;f(\#/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\mathsterling/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(\mathsterling/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\mathdollar/_{\Theta\langle F,E\rangle},f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\;\;\newline f(f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\;\;\,f(f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},f(\mathdollar,\mathdollar))/_{\Theta\langle F,E\rangle}\dot{\rightarrow}f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle}\;\;\newline f(\#/_{\Theta\langle F,E\rangle},f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(\mathsterling/_{\Theta\langle F,E\rangle},f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;\newline f(f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\;f(f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle}\,\}$.
[…]
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathdollar\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathsterling\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
@@
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathdollar\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathsterling\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\flat\dot{\rightarrow}\flat/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},\flat/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;f(\flat/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\#\,\}$.
$C\langle F,E\rangle=\{\,\#/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle}\,\}$,
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle F,E\rangle},\;\mathsterling\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\flat\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\#/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\newline f(\mathdollar/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\mathdollar/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\mathdollar/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;\newline f(\mathsterling/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\mathsterling/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle},\;f(\mathsterling/_{\Theta\langle F,E\rangle},\mathsterling/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\mathsterling/_{\Theta\langle F,E\rangle}\,\}$.
@@
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathdollar\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathsterling\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\flat\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\newline f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.
@@
$C\langle E,F\rangle=\{\,\#/_{\Theta\langle E,F\rangle},\;\mathdollar/_{\Theta\langle E,F\rangle}\,\}$.
$R\langle E,F\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E,F\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\newline f(\#/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;f(\#/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;\newline f(\mathdollar/_{\Theta\langle E,F\rangle},\#/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle},\;f(\mathdollar/_{\Theta\langle E,F\rangle},\mathdollar/_{\Theta\langle E,F\rangle})\dot{\rightarrow}\mathdollar/_{\Theta\langle E,F\rangle}\;\}$.
@@
[…]
$R\langle F,E\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},\;\mathdollar\dot{\rightarrow}\mathdollar/_{\Theta\langle F,E\rangle},\;\newline f(\#/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}\#/_{\Theta\langle F,E\rangle},f(\#/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\#,\mathdollar)/_{\Theta\langle F,E\rangle},\;\newline f(\mathdollar/_{\Theta\langle F,E\rangle},\#/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar,\#)/_{\Theta\langle F,E\rangle},\;f(\mathdollar/_{\Theta\langle F,E\rangle},\mathdollar/_{\Theta\langle F,E\rangle})\dot{\rightarrow}f(\mathdollar,\mathdollar)/_{\Theta\langle F,E\rangle}\,\}$.
@@
[…]
$R\langle E\cup F,\emptyset\rangle=\{\,\#\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;\mathdollar\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle},\;f(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})\dot{\rightarrow}\#/_{\Theta\langle E\cup F,\emptyset\rangle}\,\}$.